Sato Jewellers has had a request for a special order for 10 gold bangles for the members of a wedding party. The normal selling price of a gold bangle is $358.50 and its unit product cost is $239.00, as shown below:

  Direct materials $ 115.00
  Direct labour 70.00
  Manufacturing overhead 54.00
  Unit product cost $ 239.00


Most of the manufacturing overhead is fixed and unaffected by variations in how much jewellery is produced in any given period. However, $9 of the overhead is variable, depending on the number of bangles produced. The customer would like special filigree applied to the bangles. This filigree would require additional materials costing $8 per bangle and would also require acquisition of a special tool costing $375 that would have no other use once the special order was completed. This order would have no effect on the company’s regular sales, and the order could be filled using the company’s existing capacity without affecting any other order.


Required:
a.
What effect would accepting this order have on the company’s operating income if a special price of $328.50 is offered per bangle for this order? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. Should the special order be accepted at this price?

  • Yes

  • No
Answer #1

A. Special order revenue = 10*328.50 = 3285

Cost of making = (115+70+9+8)*10 + 375

= 2395

Financial advantage = 3285 - 2395

Financial advantage = 890

B. Yes